If there are 15 entrees 11 sides and 4 beverages on a kids menu how many possible combinatioins are there?
Answer by stanbon(75887)

If there are 15 entres 11 sides and 4 beverages on a kids menu how many possible combinations are there?
-----
Ans: 15*11*4 = 165*4 = 660 combinations.
===================
Cheers,
Stan H.
-------------------